# Reinforcement Learning Teachers of Test Time Scaling

### v1

@@ 5 Related work
Inspired by the unprecedented abilities of the OpenAI o1 model (Jaech et al., 2024), there has been a resurgence of RL approaches aimed at inducing a new kind of open-ended reasoning to scale test-time compute. The work from Guo et al. (2025) was another milestone in this new domain, providing a first openly detailed example of what is possible with large models and RL. Other follow-ups considered smaller LMs and ways to decrease optimization costs via approaches such as explicit task breakdown (Shen et al., 2025), exploration strategies (Hou et al., 2025), new RL objectives (Liu et al., 2025), and cold-start data scale (Xu et al., 2025). However, it is still an open question if RL on small models can go beyond cheaper supervised alternatives (Guo et al., 2025) and induce new skills beyond the pretraining corpus (Yue et al., 2025). In contrast to this work, RLTs break the traditional framework of maximizing one-hot accuracy with verifiable rewards – turning the task on its head by feeding the model the correct solution as input and avoiding RL’s inherent exploration challenge.
A large part of the recent test-time scaling literature considering smaller LMs has focused on inducing reasoning with “teacher-student” supervised distillation (Hinton et al., 2015), a widely validated technique in traditional LM development (Hui et al., 2024; Shing et al., 2025). This approach’s popularity to induce LM reasoning dates earlier than the RL paradigm, with older methods harnessing verifiers and prompting for self-improvement (Zelikman et al., 2022, 2024). By following a common structure of generation, filtering correct responses, and postprocessing them, modern RL-based distillation has seen significant advances mostly driven by more capable teachers (Labs, 2025; Li et al., 2025a) and carefully curating targeted datasets (Muennighoff et al., 2025; Ye et al., 2025). However, the effectiveness of current distillation pipelines was shown to be closely tied to the properties of the student itself (Li et al., 2025c; Xu et al., 2025), and their ability to induce actual generalization remains unclear (Chu et al., 2025). Unlike these traditional distillation pipelines, the RLT framework does not rely on verifiers for filtering, directly optimizes the teacher for downstream distillation, and does not require any post-processing, allowing direct transfer of reasoning capabilities to arbitrary tasks and even larger student models.
@@ 6 Discussion and extensions
This work introduced a new class of Reinforcement-Learned Teachers trained with a simpler dense-reward task that inputs both each problem’s question and solution, and optimizes the LM to provide instructive reasoning traces for distillation as outputs. Empirically, students trained or cold-started from the raw outputs of a 7B RLT obtain higher performance than prior distillation pipelines using orders-of-magnitude larger LMs. Furthermore, RLTs maintain their effectiveness even for distilling much larger students, and when providing reasoning traces for out-of-distribution tasks beyond their training corpus. Nonetheless, our work has only begun to study the design space of our new framework, with many exciting directions yet to be explored. One example is training RLTs and their students in tandem, allowing the teacher’s explanations to adapt to the student’s learning dynamics live. Pushing this further, the same model could even take both roles, iterating RL with our task formulation, providing access to each problem’s solution, to obtain instructive reasoning traces, and self-distillation (Zelikman et al., 2022) to revise its own explanation and learn how to solve questions from scratch – unifying the open-endedness of RL with the stability of supervised optimization.
@@ A.1 RLT training phase and reward
Fine-tuned model Qwen2.5-7B-instruct [Hui et al., 2024]
Number of training steps 250
Batch size 256
@@
Student model Qwen2.5-7B-instruct [Hui et al., 2024]
@@
Our experiments are conducted on a single compute node comprising 8 Nvidia Hopper H100 GPUs, 1.8 TB of memory, and 208 Intel Xeon Platinum 8481C CPUs. Due to the efficiency of training with our small 7B models, we note this setup is significantly less resource-intensive than prior RL and even SFT work, often relying on multi-node settings [Muennighoff et al., 2025, Li et al., 2025a]. As described in Section 4, for its new RL phase, we train our Reinforcement Learned Teachers on the set of questions and solutions selected by Li et al. [2025a] [^1] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k which is available under an Apache 2.0 License, comprising less than 17K math and coding problems. All our new teachers are trained from a Qwen2.5-7B-Instruct LM [Hui et al., 2024] also available under an Apache 2.0 License, together with the other models from the Qwen family. Before RL, we shortly fine-tune on pre-collected samples using example traces from Labs [2025] formatted using our new system prompt and tags. This short, inexpensive phase is conducted with the same distillation hyperparameters used for the 1K data subset following Muennighoff et al. [2025], detailed in Appendix B, but with double the number of epochs, and serves to quickly familiarize our teacher with the new RLT input format. We also use the output checkpoint at the end of this phase for our correlation analysis experiment conducted in Section 4.5. For the same reason and also to limit the requirements for reproducibility, our lightweight Qwen2.5-7B-instruct [Hui et al., 2024] student model used to compute the RLT reward function is initialized with the checkpoint provided by Labs [2025] [^2] https://huggingface.co/bespokelabs/Bespoke-Stratos-7B available under an Apache 2.0 License, which is already familiar with their system prompt without needing further training. However, we found that the RLT reward is robust to the specific LM choice for the student, yielding numerically close values when using our own Qwen student finetuned on only 1K samples.
@@
We use simple coefficients to regulate the terms in the RLT rewards, with $\lambda=3$ to scale $r^{\mathrm{KL}}$ and $\alpha=0.01$ to scale both the min term in $r^{\mathrm{SS}}$ and the max term in $r^{\mathrm{KL}}$. Our choice was based on making each individual component of the RLT rewards have approximately the same expected magnitude over the model’s initial output completions. As we observed the overall rankings to be quite robust after testing small changes to these initial choices, we did not find extensive sweeps necessary. We also add a -1 penalty to any completion that does not use the explanation tags or that exceeds our maximum generation context length to limit training time and disincentivize overly long and expensive reasoning traces. For our training runs, we make use of a custom GRPO implementation with the specifications from Shao et al. [2024], extending the TRL library [von Werra et al., 2020] with faster distributed VLLM generation [Kwon et al., 2023]. Our RL phase is short, comprising only 250 steps, less than a single epoch, with a small batch size of $256$, an AdamW optimizer [Loshchilov, 2017] with a constant learning rate of $1\times 10^{-6}$, and a group size of $64$. We synchronize the reference model every 32 steps as popularized by [Gorbatovski et al., 2024] with a mixup ratio of 0.9. For the post-cold-start RL phase employed in Section 4.3, we use our same GRPO implementation with the standard student format and correctness-based rewards described in Section 2. The only difference is that we train with a batch size of 1024, and for one epoch in total on the very small LIMR [Li et al., 2025b] dataset, containing less than 1K samples. We provide a full list of hyperparameters to ensure reproducibility in Table 3. All main libraries used in our implementation are again available under an Apache 2.0 Licence.
@@ A.2 Traditional reasoning and RLT formats
As detailed in Section 3.2 and illustrated in Figure 2, RLTs are prompted with a new formatting style that provides both the question and solution to each problem as inputs. Then, their system prompt instructs to output instructive, detailed step-by-step explanations, connecting the dots between the two. In contrast, traditional formats used for reasoning datasets employed for RL and distillation forego any information about each problem’s solution and task the model to solve each problem from scratch. We provide a specific comparison between the two, contrasting the system prompts used for traditional RL and distillation from Li et al. [2025a] in Figure 6, and our own new RLT input format in Figure 7. As shown, our prompting format’s design strives to make minimal changes to the prompts used for traditional reasoning frameworks, replacing surrounding explanation with think tags, and simply appending the solution tokens after the user’s provided input question, to allow our teachers to make full use of this information before generating each completion. This design also practically avoids the need to re-prompt our teacher multiple times for each question and manually check the answers to filter out incorrect solutions.
@@
Figure 6: Reasoning input format employed in traditional RL and student distillation, using an example question and the system prompt from Li et al. [2025a], providing the model instructions first to present a step-by-step rationale and then the problem’s solution, deriving it from scratch.
@@
Figure 7: RLT Input format using an example question from Li et al. [2025a], providing the model instructions to produce a step-by-step explanation given the solution to each problem as input.
@@ B.1 Reasoning traces generation
As detailed in Sections 3 and 4, we collect reasoning traces for each sample in the student distillation dataset by feeding our RLTs both the problem’s question and its solution as input. Moreover, for our cold-starting experiments in Section 4.3, we also collect reasoning traces by feeding the RL-trained Qwen models each question and postprocessing each output with GPT4.1 [Achiam et al., 2023] similarly to Li et al. [2025a]. We provide the hyperparameters used to collect these traces across all our datasets and settings in Table 4. In particular, we use standard generation hyperparameters for reasoning Qwen-based models [Li et al., 2025a], that are aligned with the online generation parameters during our new RL phases, as detailed in Appendix A. The only main difference is that we allow for a longer maximum context size to avoid collecting incomplete traces for downstream distillation, which is possible thanks to the fact that at test time we are not subject to the same training-time memory constraints dictated by backpropagating through long traces.
The RLT distillation datasets used for training all of our 7B students were collected by generating a single completion for each question-solution pair, directly placing it in the student format. After preliminary experiments, we also found our 32B students can be particularly sensitive to cropped reasoning traces that exceed the 16384 maximum context length specified in the SFT hyperparameters from Li et al. [2025a], which they purposefully limited for computational efficiency. Thus, to avoid this mismatch potentially affecting our results, we simply collected up to 16 reasoning traces for each question-solution pair and selected anyone below 16384, otherwise resorting to $r^{\mathrm{SS}}$ for selection.
Table 5: Hyperparameter listing student distillation phases on the datasets from the set of questions from Li et al. [2025a], and the countdown dataset.
@@
Distilled model Qwen2.5-7B/32B-instruct [Hui et al., 2024]
@@
Learning rate $\times 10^{-5} 1×1$$\times 10^{-5}$ $1\times 10^{-5}$
@@
Final learning rate $\times 10^{-6} 1×1$$\times 10^{-6}$ $1\times 10^{-6}$
@@
Optimizer AdamW [Loshchilov, 2017] AdamW [Loshchilov, 2017] AdamW [Loshchilov, 2017]
@@
Adam epsilon $\times 10^{-8} 1×1$$\times 10^{-8}$ $1\times 10^{-8}$
@@ B.2 Student distillation specifications
Our main experiments from Section 4 use the RLT traces collected either from the full set of training questions from Li et al. [2025a] or its randomly selected 1K data subset. For the student distillation phases of these experiments, we re-use the same hyperparameters from Li et al. [2025a] and Muennighoff et al. [2025], for the full data and 1K subset, respectively. The purpose of not re-tuning based on our new data is to attempt to isolate the reasoning traces used for student distillation as the only degree of variation in our comparison between RLTs with traditional reasoning distillation pipelines. Furthermore, in our experiments in Section 4.4, we also compare transferring students learned on our original set of datapoints and direct RL with transferring the RLTs themselves zero-shot to the countdown task [Gandhi et al., 2024]. For these experiment, we use a set of 16K automatically-generated countdown question and solution pairs with 3 or 4 numbers and, after obtaining the corresponding RLT traces, distill our students using the hyperparameters from Li et al. [2025a] once again, which we found to work well in practice without further tuning, given the two similar dataset sizes.
We provide a full list of the hyperparameters used for all our distillation experiments in each setting in Table 5, where we highlight the key differences across the three. As detailed, the considered approaches mostly differ in terms of the number of epochs, the batch size, and the optimizer parameters. These differences reflect the total number of samples and relative variance in each data batch. In line with the findings from the relative prior works [Muennighoff et al., 2025, Li et al., 2025a, Team, 2025b], we note that training on these small datasets was very inexpensive and could be completed within hours. We note that the learning rate of our reinforcement learning phases matches the final learning rate during our SFT optimization, a simple choice which we found to work well in practice.
Table 6: Hyperparameter listing for the traditional RL on the Li et al. [2025b] dataset.
@@
Fine-tuned model Qwen2.5-7B-instruct [Hui et al., 2024]/Bespokes-7B [Labs, 2025]
@@
Learning rate $\times 10^{-6} 1×1$$\times 10^{-6}$
@@
Final learning rate $\times 10^{-6} 1×1$$\times 10^{-6}$
@@
Optimizer AdamW [Loshchilov, 2017] AdamW [Loshchilov, 2017]
@@
Adam epsilon $\times 10^{-8} 1×1$$\times 10^{-8}$
@@ B.3 Student RL details
For our cold-starting experiments in Section 4.3, we also implement and perform a phase of traditional RL training optimizing the model with correctness-based rewards under the “student’s perspective.” This phase is done atop a Qwen2.5-7B-Instruct, the cold-started Bespoke-7B, and the RLT-7B models using our same GRPO implementation on the open-source LIMR dataset [Li et al., 2025b]. We re-use most of the hyperparameters from the RLT training phase, with a larger batch size of 1024 to better cope with the increased variance and reward sparsity of traditional RL. Additionally, we also conduct traditional RL as a baseline for our out-of-distribution transfer experiments in Section 4.4 that directly optimizes correctness on the countdown task starting again from the Qwen2.5-7B-Instruct model, and the cold-started Bespoke-7B models. For this particular task, we found using a larger batch size to be unnecessary, and, as done for RLT training, we again optimized the model for 250 total steps as we observed performance convergence and sporadic instabilities beyond that point. We provide a full list of the distillation hyperparameters used for our RL phases on these tasks in Table 6, where we highlight the key differences between these two settings.
Table 7: Student evaluation hyperparameter listing for the RLT generation pipeline, matching the hyperparameters from Li et al. [2025a].

### v2

@@ 5 Related work
Inspired by the unprecedented abilities of the OpenAI o1 model [33], there has been a resurgence of RL approaches aimed at inducing a new kind of open-ended reasoning to scale test-time compute. The work from Guo et al. [7] was another milestone in this new domain, providing a first openly detailed example of what is possible with large models and RL. Other follow-ups considered smaller LMs and ways to decrease optimization costs via approaches such as explicit task breakdown [10], exploration strategies [34], new RL objectives [35], and cold-start data scale [9]. However, it is still an open question if RL on small models can go beyond cheaper supervised alternatives [7] and induce new skills beyond the pretraining corpus [6]. In contrast to this work, RLTs break the traditional framework of maximizing one-hot accuracy with verifiable rewards – turning the task on its head by feeding the model the correct solution as input and avoiding RL’s inherent exploration challenge.
A large part of the recent test-time scaling literature considering smaller LMs has focused on inducing reasoning with “teacher-student” supervised distillation [36], a widely validated technique in traditional LM development [22, 37]. This approach’s popularity to induce LM reasoning dates earlier than the RL paradigm, with older methods harnessing verifiers and prompting for self-improvement [38, 39]. By following a common structure of generation, filtering correct responses, and postprocessing them, modern RL-based distillation has seen significant advances mostly driven by more capable teachers [12, 11] and carefully curating targeted datasets [8, 16]. However, the effectiveness of current distillation pipelines was shown to be closely tied to the properties of the student itself [40, 9], and their ability to induce actual generalization remains unclear [41]. Unlike these traditional distillation pipelines, the RLT framework does not rely on verifiers for filtering, directly optimizes the teacher for downstream distillation, and does not require any post-processing, allowing direct transfer of reasoning capabilities to arbitrary tasks and even larger student models.
@@ 6 Discussion and extensions
This work introduced a new class of Reinforcement-Learned Teachers trained with a simpler dense-reward task that inputs both each problem’s question and solution, and optimizes the LM to provide instructive reasoning traces for distillation as outputs. Empirically, students trained or cold-started from the raw outputs of a 7B RLT obtain higher performance than prior distillation pipelines using orders-of-magnitude larger LMs. Furthermore, RLTs maintain their effectiveness even for distilling much larger students, and when providing reasoning traces for out-of-distribution tasks beyond their training corpus. Nonetheless, our work has only begun to study the design space of our new framework, with many exciting directions yet to be explored. One example is training RLTs and their students in tandem, allowing the teacher’s explanations to adapt to the student’s learning dynamics live. Pushing this further, the same model could even take both roles, iterating RL with our task formulation, providing access to each problem’s solution, to obtain instructive reasoning traces, and self-distillation [38] to revise its own explanation and learn how to solve questions from scratch – unifying the open-endedness of RL with the stability of supervised optimization.
Acknowledgments
The authors would like to thank Takuya Akiba for insightful project discussions, together with Yutaro Yamada and Boris Meinardus for providing feedback on early versions of the text.
@@ A.1 RLT training phase and reward
Fine-tuned model Qwen2.5-7B-instruct [22]
Number of training steps 125
Batch size 1024
@@
Student model Qwen2.5-7B-instruct [22]
@@
Our experiments are conducted on a single compute node comprising 8 Nvidia Hopper H100 GPUs, 1.8 TB of memory, and 208 Intel Xeon Platinum 8481C CPUs. Due to the efficiency of training with our small 7B models, we note this setup is significantly less resource-intensive than prior RL and even SFT work, often relying on multi-node settings [8, 11]. As described in Section 4, for its new RL phase, we train our Reinforcement Learned Teachers on the set of questions and solutions selected by Li et al. [11] [^2] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k which is available under an Apache 2.0 License, comprising less than 17K math and coding problems. All our new teachers are trained from a Qwen2.5-7B-Instruct LM [22] also available under an Apache 2.0 License, together with the other models from the Qwen family. Before RL, we shortly fine-tune on pre-collected samples using example traces from Labs [12] formatted using our new system prompt and tags. This short, inexpensive phase is conducted with the same distillation hyperparameters used for the 1K data subset following Muennighoff et al. [8], detailed in Appendix B, but with double the number of epochs, and serves to quickly familiarize our teacher with the new RLT input format. We also use the output checkpoint at the end of this phase for our correlation analysis experiment conducted in Section 4.5. For the same reason and also to limit the requirements for reproducibility, our lightweight Qwen2.5-7B-instruct [22] student model used to compute the RLT reward function is initialized with the checkpoint provided by Labs [12] [^3] https://huggingface.co/bespokelabs/Bespoke-Stratos-7B available under an Apache 2.0 License, which is already familiar with their system prompt without needing further training. However, we found that the RLT reward is robust to the specific LM choice for the student, yielding numerically close values when using our own Qwen student finetuned on only 1K samples.
@@
We use simple coefficients to regulate the terms in the RLT rewards, with $\lambda=3$ to scale $r^{\mathrm{KL}}$ and $\alpha=0.01$ to scale both the min term in $r^{\mathrm{SS}}$ and the max term in $r^{\mathrm{KL}}$. Our choice was based on making each individual component of the RLT rewards have approximately the same expected magnitude over the model’s initial output completions. As we observed the overall rankings to be quite robust after testing small changes to these initial choices, we did not find extensive sweeps necessary. We also add a -1 penalty to any completion that does not use the explanation tags or that exceeds our maximum generation context length to limit training time and disincentivize overly long and expensive reasoning traces. For our training runs, we make use of a custom GRPO implementation with the specifications from Shao et al. [14], extending the TRL library [43] with faster distributed VLLM generation [44]. Our RL phase is short, comprising only 125 steps, less than a single epoch, with a batch size of $1024$, an AdamW optimizer [42] with a constant learning rate of $1\times 10^{-6}$, and a group size of $64$. We synchronize the reference model every 32 steps as popularized by [45] with a mixup ratio of 0.9. For the post-cold-start RL phase employed in Section 4.3, we use our same GRPO implementation with the standard student format and correctness-based rewards described in Section 2. The only difference is that we train for one epoch in total on the very small LIMR [29] dataset, containing less than 1K samples. We provide a full list of hyperparameters to ensure reproducibility in Table 3. All main libraries used in our implementation are again available under an Apache 2.0 Licence.
@@ A.2 Traditional reasoning and RLT formats
As detailed in Section 3.2 and illustrated in Figure 2, RLTs are prompted with a new formatting style that provides both the question and solution to each problem as inputs. Then, their system prompt instructs to output instructive, detailed step-by-step explanations, connecting the dots between the two. In contrast, traditional formats used for reasoning datasets employed for RL and distillation forego any information about each problem’s solution and task the model to solve each problem from scratch. We provide a specific comparison between the two, contrasting the system prompts used for traditional RL and distillation from Li et al. [11] in Figure 6, and our own new RLT input format in Figure 7. As shown, our prompting format’s design strives to make minimal changes to the prompts used for traditional reasoning frameworks, replacing surrounding explanation with think tags, and simply appending the solution tokens after the user’s provided input question, to allow our teachers to make full use of this information before generating each completion. This design also practically avoids the need to re-prompt our teacher multiple times for each question and manually check the answers to filter out incorrect solutions.
@@
Figure 6: Reasoning input format employed in traditional RL and student distillation, using an example question and the system prompt from Li et al. [11], providing the model instructions first to present a step-by-step rationale and then the problem’s solution, deriving it from scratch.
@@
Figure 7: RLT Input format using an example question from Li et al. [11], providing the model instructions to produce a step-by-step explanation given the solution to each problem as input.
@@ B.1 Reasoning traces generation
As detailed in Sections 3 and 4, we collect reasoning traces for each sample in the student distillation dataset by feeding our RLTs both the problem’s question and its solution as input. Moreover, for our cold-starting experiments in Section 4.3, we also collect reasoning traces by feeding the RL-trained Qwen models each question and postprocessing each output with GPT4.1 [30] similarly to Li et al. [11]. We provide the hyperparameters used to collect these traces across all our datasets and settings in Table 4. In particular, we use standard generation hyperparameters for reasoning Qwen-based models [11], that are aligned with the online generation parameters during our new RL phases, as detailed in Appendix A. The only main difference is that we allow for a longer maximum context size to avoid collecting incomplete traces for downstream distillation, which is possible thanks to the fact that at test time we are not subject to the same training-time memory constraints dictated by backpropagating through long traces.
The RLT distillation datasets used for training all of our 7B students were collected by generating a single completion for each question-solution pair, directly placing it in the student format. After preliminary experiments, we also found our 32B students can be particularly sensitive to cropped reasoning traces that exceed the 16384 maximum context length specified in the SFT hyperparameters from Li et al. [11], which they purposefully limited for computational efficiency. Thus, to avoid this mismatch potentially affecting our results, we simply collected up to 16 reasoning traces for each question-solution pair and selected anyone below 16384, otherwise resorting to $r^{\mathrm{SS}}$ for selection.
Table 5: Hyperparameter listing student distillation phases on the datasets from the set of questions from Li et al. [11], and the countdown dataset.
@@
Distilled model Qwen2.5-7B/32B-instruct [22]
@@
Learning rate $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
@@
Final learning rate $1\times 10^{-6}$ $1\times 10^{-6}$ $1\times 10^{-6}$
@@
Optimizer AdamW [42] AdamW [42] AdamW [42]
@@
Adam epsilon $1\times 10^{-8}$ $1\times 10^{-8}$ $1\times 10^{-8}$
@@ B.2 Student distillation specifications
Our main experiments from Section 4 use the RLT traces collected either from the full set of training questions from Li et al. [11] or its randomly selected 1K data subset. For the student distillation phases of these experiments, we re-use the same hyperparameters from Li et al. [11] and Muennighoff et al. [8], for the full data and 1K subset, respectively. The purpose of not re-tuning based on our new data is to attempt to isolate the reasoning traces used for student distillation as the only degree of variation in our comparison between RLTs with traditional reasoning distillation pipelines. Furthermore, in our experiments in Section 4.4, we also compare transferring students learned on our original set of datapoints and direct RL with transferring the RLTs themselves zero-shot to the countdown task [32]. For these experiment, we use a set of 16K automatically-generated countdown question and solution pairs with 3 or 4 numbers and, after obtaining the corresponding RLT traces, distill our students using the hyperparameters from Li et al. [11] once again, which we found to work well in practice without further tuning, given the two similar dataset sizes.
We provide a full list of the hyperparameters used for all our distillation experiments in each setting in Table 5, where we highlight the key differences across the three. As detailed, the considered approaches mostly differ in terms of the number of epochs, the batch size, and the optimizer parameters. These differences reflect the total number of samples and relative variance in each data batch. In line with the findings from the relative prior works [8, 11, 31], we note that training on these small datasets was very inexpensive and could be completed within hours. We note that the learning rate of our reinforcement learning phases matches the final learning rate during our SFT optimization, a simple choice which we found to work well in practice.
Table 6: Hyperparameter listing for the traditional RL on the Li et al. [29] dataset.
@@
Fine-tuned model Qwen2.5-7B-instruct [22]/Bespokes-7B [12]
@@
Learning rate $1\times 10^{-6}$ $1\times 10^{-6}$
@@
Final learning rate $1\times 10^{-6}$ $1\times 10^{-6}$
@@
Optimizer AdamW [42] AdamW [42]
@@
Adam epsilon $1\times 10^{-8}$ $1\times 10^{-8}$
@@ B.3 Student RL details
For our cold-starting experiments in Section 4.3, we also implement and perform a phase of traditional RL training optimizing the model with correctness-based rewards under the “student’s perspective.” This phase is done atop a Qwen2.5-7B-Instruct, the cold-started Bespoke-7B, and the RLT-7B models using our same GRPO implementation on the open-source LIMR dataset [29]. We re-use most of the hyperparameters from the RLT training phase with a batch size of 1024, something particularly important to cope with the increased variance and reward sparsity of traditional RL. Additionally, we also conduct traditional RL as a baseline for our out-of-distribution transfer experiments in Section 4.4 that directly optimizes correctness on the countdown task starting again from the Qwen2.5-7B-Instruct model, and the cold-started Bespoke-7B models. For this particular task, we found using a larger batch size to be not strictly necessary, as for RLT training, and obtained better results optimizing the model for more steps (250 total) with a batch size of $256$. We provide a full list of the distillation hyperparameters used for our RL phases on these tasks in Table 6, where we highlight the key differences between these two settings.
Table 7: Student evaluation hyperparameter listing for the RLT generation pipeline, matching the hyperparameters from Li et al. [11].